You can invest in an account that pays simple interest or an account that pays compound interest. In either case, you plan to invest $3,800 today and both accounts have an annual interest rate of 6 percent. How much more interest will you receive in the 10th year in the account that pays compound interest?
Answer #1

Value of Investment in Simple Interest = 3,800(1 + 10(0.06))

Value of Investment in Simple Interest = $6,080

Value in Investment in Compound Interest = 3,800(1.06)10

Value in Investment in Compound Interest = $6,805.22

Difference in Interest = 6,805.22 - 6,080

Difference in Interest = $725.22